Elimination
4x-6y=3
6x-4y=-3

Answer by stanbon(75887)

4x-6y=3
6x-4y=-3
----
Multiply thru 1st equation by 3
Multiply thru 2nd equation by 2
-----
12x - 18y = 9
12x - 8y = -6
----------
Substract and solve for "y":
10y = -15
y = -3/2
----
Solve for "x":
4x - 6y = 3
4x - 6(-3/2) = 3
4x + 9 = 3
4x = -6
x = -3/2
=====
Solution: x = -3/2 ; y = -3/2
==============================
cheers,
Stan H.